OpenStudy (anonymous):

Evaluate the indefinite integral of x^(3) (x^2+49)^1/2 dx

OpenStudy (anonymous):

Will most likely need to do a trig substitution. Try letting:\[x=\tan\theta\]

OpenStudy (anonymous):

I did u-substitution and let u^2 = x^2+49

OpenStudy (anonymous):

I'm not quite sure if that will lead to a solution.

OpenStudy (anonymous):

well I don't know what I'm doing haha

OpenStudy (raden):

integral by u-sub will works u = x^2 + 49 --------> x^2 = u - 49 du = 2x dx 1/2 du = x dx now see ur integral becomes int x^3 (x^2 + 49)^1/2 dx = int x x^2 (x^2 + 49)^1/2 dx = 1/2 int (u-49) u^1/2 du = 1/2 int (u^(3/2) - 49u^1/2) du

OpenStudy (raden):

that should make easier for u, now

OpenStudy (anonymous):

now do you plug x^2+49 back in for u?

OpenStudy (raden):

after integration, yes :)

OpenStudy (anonymous):

so now I'll have to take the anti derivative of u^3/2 and 49u^1/2?

OpenStudy (raden):

yup

OpenStudy (anonymous):

I got 3/5u^5/2 and 98/3u^3/2

OpenStudy (raden):

that's 3/5 or 2/5 ?

OpenStudy (anonymous):

whoops 2/5 i meant

OpenStudy (raden):

ok, u are right dont forget multiply by 1/2 and subtitute back u = ... endly + c

OpenStudy (anonymous):

would it be: 1/2[2/5(x^2+49)^3/2 - 49(x^2+49)1/2] +C?

OpenStudy (raden):

1/2 [2/5(x^2+49)^5/2 - 98/3(x^2+49)^3/2] +C = 1/5(x^2+49)^5/2 - 49/3(x^2+49)^3/2 +C

OpenStudy (anonymous):

is that with the 1/2 multiplied through?

OpenStudy (raden):

look my coments before above = 1/2 int (u-49) u^1/2 du = 1/2 int (u^(3/2) - 49u^1/2) du 1/2 * ......... yeah, after integration we have to multiply that by 1/2

OpenStudy (anonymous):

Can it be simplified further after multiplying by 1/2?

OpenStudy (raden):

yeah, it have simplied above 1/2 [2/5(x^2+49)^5/2 - 98/3(x^2+49)^3/2] +C = 1/5(x^2+49)^5/2 - 49/3(x^2+49)^3/2 +C

OpenStudy (anonymous):

Awesome!! thank you for helping me with the problem!